### v1

# GTPC-SSCD: Gate-guided Two-level Perturbation Consistency-based Semi-Supervised Change Detection thanks: This work was supported in part by the Scientific Research Program of Tianjin Municipal Education Commission under Grant 2023KJ232. (Corresponding author: Rui Huang.)

Abstract
Semi-supervised change detection (SSCD) employs partially labeled data and a substantial amount of unlabeled data to identify differences between images captured in the same geographic area but at different times. However, existing consistency regularization-based SSCD methods only implement perturbations at a single level and can not exploit the full potential of unlabeled data. In this paper, we introduce a novel Gate-guided Two-level Perturbation Consistency regularization-based SSCD method (GTPC-SSCD), which simultaneously maintains strong-to-weak consistency at the image level and perturbation consistency at the feature level, thus effectively utilizing the unlabeled data. Moreover, a gate module is designed to evaluate the training complexity of different samples and determine the necessity of performing feature perturbations on each sample. This differential treatment enables the network to more effectively explore the potential of unlabeled data. Extensive experiments conducted on six public remote sensing change detection datasets demonstrate the superiority of our method over seven state-of-the-art SSCD methods.
Index Terms:
Change detection (CD), consistency regularization, remote sensing, gate machine, semi-supervised learning
I Introduction
Semi-supervised change detection (SSCD) aims to identify pixel-level changes between two images taken from the same scene at different times, using a limited amount of labeled data and a large amount of unlabeled data  [1, 2, 3]. It has wide applications in various fields, including natural resource monitoring and utilization  [4, 5], disaster monitoring and assessment [6], urban management and development [7, 8].
Semi-supervised methods can be divided into three categories: adversarial learning-based methods, pseudo-label-based methods, and consistency regularization-based methods. AdvEnt [9], SemiCDNet [10], and SALCD [11] are typical adversarial learning-based methods that employ alternative optimization strategies to improve the representation learning of their respective models. Pseudo-label-based methods, such as RC-CD [12], SemiSiROC [13], and DCF-CFe [14], emphasize the improvement of pseudo-label quality and apply some techniques to enhance feature distinctiveness. Consistency regularization-based methods, including SemiCD [15], SemiSANet [16], Semi-LCD [17], SemiBuildingChange [18], and SemiPTCD [19], enforce that perturbed images or features produce identical outputs as the original inputs [20].
Recently, the consistency regularization-based framework have been increasingly drawn to SSCD methods for its simplicity and stability. However, current methods often modify network architectures or introduce additional information based on image-level weak-to-strong consistency. For example, SemiSANet [16] constructs a siamese nested UNet with graph attention, while MTCNet [21] adds extra segmentation labels to change labels. ST-RCL [22] combines self-training and consistency learning for SSCD. Concurrently, feature-level consistency regularization remains relatively simplistic and underdeveloped, such as UniMatch’s implementation of a dropout to enforce feature consistency [23]. Thus, there exists significant potential in integrating image-level and feature-level consistency regularization to further enhance the performance and robustness of SSCD.
Figure 1: Motivation: Comparison of SSCD with different perturbations by $5\%$ labeled training data. Sup-only denotes that our method only be trained by labeled training data. Feature: Feature-level perturbation consistency method. Image: Image-level strong-to-weak consistency method. Different colors are used for a better view, i.e., white for true positive, black for true negative, red for false positive, and green for false negative.
Our research shows that employing a simple SSCD architecture and applying perturbations either at the image or feature level does not yield satisfactory performance. Fig. 1 presents the results of SSCD networks employing perturbations at different perspectives with a ratio of $5\%$ labeled data and $95\%$ unlabeled data on two CD datasets [24, 10]. Although perturbations contribute to enhancing model robustness, relying solely on image perturbations may affect pixel integrity and segmentation, and exclusive dependence on feature perturbations might struggle with image diversity.
In this paper, we explore consistency regularization at a broader scope and propose a Gate-guided Two-level Perturbation Consistency regularization-based SSCD method (GTPC-SSCD). By learning both image-level strong-to-weak consistency and feature-level perturbation consistency, our method can benefit from a more thorough comprehension of the data, leading to improved robustness and discriminability. Furthermore, performing the same treatment on all samples constrains the potential effectiveness of perturbation, particularly evident in remote sensing change detection scenarios. To tackle this issue, we develop a gate module to distinguish the difficulty of training samples and determine the application of feature perturbation accordingly. As presented in Fig. 1, the performance of our method is superior to other three methods. Our contributions are summarized as follows:
  • We separately analyze image and feature perturbation implementations in SSCD, and meticulously combine them with optimized strategies to improve performance.
  • We introduce a gate-guided two-level perturbation consistency regularization-based SSCD method, utilizing a gate mechanism to assess the difficulty of samples and decide whether to apply feature perturbations.
  • We conduct extensive experiments on six public CD datasets. Our method exhibits higher accuracy compared to other seven state-of-the-art SSCD methods.
II Methodology
Semi-supervised change detection (SSCD) utilizes a limited amount of labeled data and a large amount of unlabeled data to train a change detection network for accurate change map generation. The training set consists of two subsets, a labeled set and an unlabeled set. The labeled set can be represented as $\mathcal{D}_{l}=\{{(\mathbf{X}^{l}_{A,i},\mathbf{X}^{l}_{\mathrm{B},i}),\mathbf{Y}^{l}_{i}}\}^{M}_{i=1}$, where $(\mathbf{X}^{l}_{A,i},\mathbf{X}^{l}_{B,i})$ denotes the $i$-th labeled image pair, $\mathbf{X}^{l}_{A,i}$ is the pre-change image, $\mathbf{X}^{l}_{B,i}$ is the post-change image, and $\mathbf{Y}^{l}_{i}$ is the corresponding ground truth. Let $\mathcal{D}_{u}=\{{(\mathbf{X}^{u}_{A,i},\mathbf{X}^{u}_{B,i})}\}^{N}_{i=1}$ denotes the unlabeled set. $(\mathbf{X}^{u}_{A,i},\mathbf{X}^{u}_{B,i})$ is the $i$-th unlabeled image pair. $M$ and $N$ indicate the number of labeled and unlabeled image pairs, respectively. In most cases, we have $N>>M$. In the following sections, we will introduce the implementation details of the proposed change detection network.
Figure 2: Framework of the proposed GTPC-SSCD method.
II-A Our consistency regularization-based SSCD method
Our SSCD method consists of a supervised training part and an unsupervised training part.
In the supervised training part, we use the labeled set $\mathcal{D}_{l}$ to train the CD network $\phi$. The network processes weakly augmented image pairs to generate change map $\mathbf{P}^{l}$. We apply standard cross-entropy (CE) loss as supervision. The loss of the supervised training part is:
$\mathcal{L}_{s}=\mathcal{L}_{CE}(\mathbf{P}^{l},\mathbf{Y}^{l}).$ (1)
The unsupervised part is illustrated in Fig. 2, we apply two-level perturbations to train $\phi$ on the unlabeled set $\mathcal{D}_{u}$. Specifically, we exploit strong-to-weak consistency at the image level and perturbation consistency at the feature level. The former enhances the model’s robustness to various data transformations, while the latter improves the model’s generalization ability and resistance to interference.
Strong-to-weak consistency learning. Drawing inspiration from ConMatch [25] and ReMixMatch [26], we increase the number of strong augmentation branches and ensure the strong-to-weak consistency, fully maximizing the benefits of strong augmentation. Specifically, we apply two independent strong augmentations on $(\mathbf{X}^{uw}_{A},\mathbf{X}^{uw}_{B})$, resulting in $(\mathbf{X}^{us_{1}}_{A},\mathbf{X}^{us_{1}}_{B})$ and $(\mathbf{X}^{us_{2}}_{A},\mathbf{X}^{us_{2}}_{B})$. They are fed into $\phi$ and generate three change maps $\mathbf{P}^{uw}$, $\mathbf{P}^{us_{1}}$ and $\mathbf{P}^{us_{2}}$.
We use $\mathbf{P}^{uw}$ to generate pseudo-label $\hat{\mathbf{Y}}^{uw}$ by:
$\hat{\mathbf{Y}}^{uw}=\begin{cases}1,&if\ \mathbf{P}^{uw}>\tau\\ 0,&else\end{cases}\vspace{-2pt}$ (2)
where $\tau$ = 0.95 is a confidence threshold. The image-level consistency loss of the unsupervised part is as:
$\mathcal{L}_{ui}=\frac{1}{2}(\mathcal{L}_{CE}(\mathbf{P}^{us_{1}},\hat{\mathbf{Y}}^{uw})+\mathcal{L}_{CE}(\mathbf{P}^{us_{2}},\hat{\mathbf{Y}}^{uw})).\vspace{-2pt}$ (3)
In our experiments, weak augmentations consist of random resizing, cropping, and random horizontal flipping. Strong augmentations include random color jittering, Gaussian blur, and CutMix [27].
Perturbation consistency learning. We introduce feature-level consistency by applying several different perturbations on change features and enforcing the perturbation consistency. The process of generating change maps $\mathbf{P}^{uf}_{1:K}$ is depicted as:
$\mathbf{P}^{uf}_{1:K}=f^{{}^{\prime}}_{d,1:K}(\mathbf{D}^{u}_{4},\mathbf{D}^{u^{\prime}}_{1,1:K}),\vspace{-2pt}$ (4)
where $\mathbf{D}^{u^{\prime}}_{1,1:K}=\{\mathbf{D}^{u^{\prime}}_{1,k}\}^{K}_{k=1}$, represents the difference features after perturbation. $f^{{}^{\prime}}_{d,1:K}=\{f^{{}^{\prime}}_{d,k}\}^{K}_{k=1}$, denotes the auxiliary decoders. $\mathbf{P}^{uf}_{1:K}=\{\mathbf{P}^{uf}_{k}\}^{K}_{k=1}$. The feature-level consistency loss $\mathcal{L}_{uf}$ is expressed as:
$\mathcal{L}_{uf}=\frac{1}{K}\sum^{K}_{k=1}\mathcal{L}_{CE}(\mathbf{P}^{uf}_{k},\hat{\mathbf{Y}}^{uw}),\vspace{-2pt}$ (5)
where $K$ is the total number of the auxiliary decoders. We apply seven types of feature perturbations [28], i.e., feature noise, feature dropout, object masking, context masking, guided cutout, Intermediate VAT [29], and random dropout.
Hardness analysis based gate mechanism. We add an additional decoder to produce $\mathbf{P}^{uw^{\prime}}_{i}$ and conduct an Intersection over Union (IoU) comparison with $\mathbf{P}^{uw}_{i}$ to assess the complexity of each sample. Utilizing a gate mechanism, we select challenging samples and apply various perturbations on $\mathbf{D}^{u}_{1}$ of these selected samples to obtain perturbed features. The process can be formulated as:
$iou_{i}=\mathrm{IoU}(\mathbf{P}^{uw^{\prime}}_{i},\mathbf{P}^{uw}_{i}),\vspace{-1pt}$ (6)
$\mathbf{D}^{u^{\prime}}_{1,1:K}=\begin{cases}\mathrm{FP}_{1:K}(\mathbf{D}^{u}_{1}),&if\ iou_{i}\geq\mathrm{Mid}(iou)\\ \mathbf{D}^{u}_{1},&else\end{cases}$ (7)
where $iou_{i}$ represents the IoU score of the $i$-th sample, $\mathrm{Mid}(iou)$ is the median IoU score. $\mathrm{FP}_{1:K}(\cdot)$ denotes $K$ different perturbations to the input features, respectively.
Loss function. The total loss consists of the supervised loss $\mathcal{L}_{s}$, the image-level consistency loss $\mathcal{L}_{ui}$ and the feature-level consistency loss $\mathcal{L}_{uf}$. It can be expressed as:
$\mathcal{L}=\lambda_{1}\mathcal{L}_{s}+\lambda_{2}\mathcal{L}_{ui}+\lambda_{3}\mathcal{L}_{uf},$ (8)
where $\lambda_{1}=0.5$, $\lambda_{2}=0.25$, and $\lambda_{3}=0.25$.
II-B Implementation detail
II-B1 Change Detection Network
Our CD network $\phi$ consists of two components: a difference feature generator $f_{g}$ and a decoder $f_{d}$. The process of generating change map $\mathbf{P}$ is simply expressed by:
$\mathbf{P}=\phi(\mathbf{X}_{A},\mathbf{X}_{B})=f_{d}(f_{g}(\mathbf{X}_{A},\mathbf{X}_{B})).$ (9)
Difference feature generator. The feature encoder is built on ResNet50 [30] with a siamese setup. Deep features contain rich semantic information, while shallow features encompass abundant details. We use the features of the first and fourth residual modules to calculate the difference features $\mathbf{D}_{i}$ by:
$\mathbf{D}_{i}=|\mathbf{C}^{A}_{i}-\mathbf{C}^{B}_{i}|,i=1,4,$ (10)
where $\mathbf{C}^{A}_{i}$ and $\mathbf{C}^{B}_{i}$ are the features of the $i$-th residual module from $\mathbf{X}_{A}$ and $\mathbf{X}_{B}$, respectively. ||$\cdot$|| ⋅ | is the absolute operation.
Decoder. The decoder is used to process change features and generate change maps. We apply an Atrous Spatial Pyramid Pooling (ASPP) [31] on $\mathbf{D}_{4}$ to obtain richer information $\mathbf{F}_{4}$. Next, we combine $\mathbf{D}_{1}$ and the upsampled feature of $\mathbf{F}_{4}$ and employ a classifier to generate $\mathbf{P}$. The process can be formulated as:
$\mathbf{F}_{4}$ $=\mathrm{ASPP}(\mathbf{D}_{4}),$ (11)
$\mathbf{F}_{1}=\mathrm{CBR}_{3}(\mathrm{CBR}_{3}$ $([\mathrm{Up(\mathbf{F}_{4})},\mathrm{CBR}_{1}(\mathbf{D}_{1})])),$ (12)
$\mathbf{P}$ $=\mathrm{Conv}_{1}(\mathbf{F}_{1}),$ (13)
where $\mathrm{ASPP(\cdot)}$ denotes the ASPP process, $\mathrm{Up}(\cdot)$ is upsampling operation, [,][$\cdot$,$\cdot$][ ⋅ , ⋅ ] denotes concatenate operation. $\mathrm{CBR}_{k}(\cdot)$ denotes a $k\times k$ convolutional layer followed with Batch Normalization and ReLU, and $\mathrm{Conv}_{1}(\cdot)$ represents a $1\times 1$ convolutional layer.
II-B2 Super-parameters
We use PyTorch to conduct experiments on an NVIDIA RTX2080Ti GPU. Our model adopts the SGD optimizer with learning rate of 0.02, momentum of 0.9, and weight decay of 1e-4, respectively. All models are trained for 80 epochs with batch size of 4 for both labeled and unlabeled data.
III Experiments
III-A Experimental Setups
Baselines. We compare the proposed method with seven existing state-of-the-art methods, including AdvEnt [9], s4GAN [32], SemiCDNet [10], SemiCD [15], RC-CD [12], SemiPTCD [19], and UniMatch [23]. Sup-only refers to our method that exclusively utilizes labeled data for training.
Datasets. We conduct experiments on six widely used remote sensing CD datasets, namely WHU-CD [24], LEVIR-CD [7], BCD [24], GZ-CD [10], EGY-BCD [33], and CDD [34]. Table I details our datasets, including number of image pairs, image size, train/val/test splits, spatial resolution, and change types. The first five CD datasets mainly focus on building changes, while the last CD dataset consists of multiple types of changes, including buildings, vehicles, road expansions, and more. All images are cropped into non-overlapping patches of size $256\times 256$, which are then divided into training, validation, and test sets. The training set is further divided into labeled and unlabeled data with the following ratios: $[5\%,95\%]$, $[10\%,90\%]$, $[20\%,80\%]$, $[40\%,60\%]$.
TABLE I: Dataset details.
Dataset Image pairs Image size Train/Val/Test Spatial resolution Changes
WHU-CD 1 $15354\times 32507\times 3$ $5974/743/744$ 0.2 m/pixel building
LEVIR-CD 637 $1024\times 1024\times 3$ $7120/1024/2048$ 0.5 m/pixel building
BCD 1922 $256\times 256\times 3$ $1538/192/192$ 0.2 m/pixel building
GZ-CD 19 $1006\times 1168\times 3,$ $4936\times 5224\times 3$ $2882/360/361$ 0.55 m/pixel building
EGY-BCD 6091 $256\times 256\times 3$ $4264/1218/609$ 0.25 m/pixel building
CDD 16000 $256\times 256\times 3$ $10000/3000/3000$ 0.03-1 m/pixel building, car, tree, road, etc.
Criterion. Following Bandara et al. [15], Mao et al. [19], and Yang et al. [23], we adopt intersection over union (IoU) and overall accuracy (OA) as the primary performance metric to evaluate different change detection methods.
TABLE II: Quantitative comparison of different methods on six CD datasets. The highest scores are marked in bold.
Method WHU-CD LEVIR-CD
5% 10% 20% 40% 5% 10% 20% 40%
IoU OA IoU OA IoU OA IoU OA IoU OA IoU OA IoU OA IoU OA
AdvEnt [9] 57.7 97.87 60.5 97.79 69.5 98.50 76.0 98.91 67.1 98.15 70.8 98.38 74.3 98.59 75.9 98.67
s4GAN [32] 57.3 97.94 58.0 97.81 67.0 98.41 74.3 98.85 66.6 98.16 72.2 98.48 75.1 98.63 76.2 98.68
SemiCDNet [10] 56.2 97.78 60.3 98.02 69.1 98.47 70.5 98.59 67.4 98.11 71.5 98.42 74.9 98.58 75.5 98.63
SemiCD [15] 65.8 98.37 68.0 98.45 74.6 98.83 78.0 99.01 74.2 98.59 77.1 98.74 77.9 98.79 79.0 98.84
RC-CD [12] 58.0 98.01 61.7 98.00 74.0 98.83 73.9 98.85 74.0 98.52 76.1 98.65 77.1 98.70 77.6 98.72
SemiPTCD [19] 74.1 98.85 74.2 98.86 76.9 98.95 80.8 99.17 71.2 98.39 75.9 98.65 76.6 98.65 77.2 98.74
UniMatch [23] 78.7 99.11 79.6 99.11 81.2 99.18 83.7 99.29 82.1 99.03 82.8 99.07 82.9 99.07 83.0 99.08
Sup-only 41.4 96.36 55.0 97.28 49.3 96.98 65.6 98.26 71.8 98.46 78.0 98.78 77.7 98.76 78.9 98.83
Ours 83.0 99.30 82.8 99.28 84.2 99.31 85.4 99.37 83.2 99.09 83.1 99.08 83.7 99.12 83.5 99.10
BCD GZ-CD
AdvEnt [9] 62.4 91.93 69.1 93.29 75.2 94.85 77.1 95.14 56.7 95.52 57.5 95.99 70.3 97.28 70.8 97.29
s4GAN [32] 57.7 90.77 70.9 93.95 74.9 94.64 78.0 95.35 59.4 96.13 61.6 96.23 68.5 97.10 69.4 97.08
SemiCDNet [10] 59.4 91.49 70.3 93.83 74.4 94.65 77.7 95.33 57.9 95.38 54.9 95.52 68.9 97.16 69.7 97.20
SemiCD [15] 60.9 92.08 73.1 94.23 77.5 95.26 79.3 95.66 59.5 96.27 58.6 96.03 67.0 97.03 71.5 97.36
RC-CD [12] 75.4 94.79 77.7 95.18 79.7 95.77 80.6 95.95 62.2 96.26 63.9 96.55 74.1 97.69 74.2 97.57
UniMatch [23] 78.3 96.17 79.9 96.48 80.7 96.58 81.7 96.79 68.7 97.06 69.5 97.41 72.8 97.71 71.1 97.48
Sup-only 59.3 91.49 63.3 92.48 74.1 95.16 70.1 95.16 48.9 94.49 36.9 91.94 53.5 95.44 54.4 95.73
Ours 79.9 96.52 81.3 96.76 81.3 96.78 82.4 96.95 73.9 97.71 70.7 97.36 77.8 98.09 78.4 98.17
EGY-BCD CDD
AdvEnt [9] 52.0 95.30 58.1 96.26 59.8 96.46 63.8 96.94 63.8 94.98 72.7 96.32 79.0 97.22 82.8 97.72
s4GAN [32] 53.2 95.62 56.5 96.26 59.4 96.62 64.1 96.83 62.5 94.92 70.4 96.12 78.7 97.20 82.8 97.72
SemiCDNet [10] 52.7 95.36 56.9 96.02 59.8 96.53 63.6 96.96 64.3 95.01 72.5 95.88 79.1 97.23 82.6 97.73
SemiCD [15] 54.3 95.79 59.2 96.29 61.8 96.61 65.4 96.96 66.4 95.39 74.9 96.71 81.2 97.54 84.4 97.93
RC-CD [12] 59.0 96.17 61.6 96.51 64.6 96.79 67.7 97.09 69.2 95.82 73.4 96.29 80.1 97.32 82.7 97.67
UniMatch [23] 62.8 96.74 65.5 97.10 63.6 96.91 67.3 97.26 74.8 96.69 80.3 97.41 86.8 98.30 90.7 98.81
Sup-only 44.5 94.03 54.9 95.65 44.8 94.28 60.0 96.37 49.7 92.58 62.3 94.71 66.2 95.32 78.5 97.15
Ours 64.5 96.91 65.9 97.15 66.1 97.13 68.9 97.44 76.4 96.89 82.3 97.70 88.0 98.46 90.2 98.74
Figure 3: Detection results of different methods on six CD datasets at the $5\%$ labeled training ratio. Different colors are used for a better view, i.e., white for true positive, black for true negative, red for false positive, and green for false negative.
III-B Results and Discussion
Comparison to the State-of-the-Art. Table II exhibits the quantitative results of different methods on WHU-CD, LEVIR-CD, BCD, GZ-CD, EGY-BCD, and CDD datasets. From the findings in these tables, we can draw the following observations. ❶ Most SSCD methods outperform Sup-only under the same partition, confirming the benefit of unlabeled data for SSCD. ❷ Our method achieves the best performance across all four partitions, notably surpassing other methods, especially on WHU-CD and GZ-CD. On WHU-CD, compared to the current SOTA method UniMatch, our method brings $4.3\%$, $3.2\%$, $3.0\%$, and $1.7\%$ performance gain in terms of IoU with $5\%$, $10\%$, $20\%$, and $40\%$ labeled training data, respectively. On GZ-CD, the improved performance with IoU of our method over the best UniMatch are $5.2\%$, $1.2\%$, $5.0\%$, and $7.3\%$.
Detection results. Fig. 3 shows some examples of different methods on six CD datasets under the partition of $5\%$. Our method exhibits higher accuracy and richer details. Both quantitative and qualitative results support our method’s superiority.
Computational complexity analysis. Table III shows a detailed comparison of parameters and computation costs of different methods. Compared to other methods, our method achieves a high FPS with a moderate parameter count, striking an effective balance between performance and efficiency.
TABLE III: Comparison of parameters and computational costs of different methods.
Method GFLOPs Params (M) FPS
AdvEnt 74.06 49.6 40.82
s4GAN 76.72 49.6 41.54
SemiCDNet 75.69 52.4 40.87
SemiCD 75.37 50.7 39.30
UniMatch 38.30 40.5 52.19
Ours 65.87 57.3 51.64
Ablation study. Table IV illustrates four sets of ablation studies. From this table, we have the following findings. ❶ Under the “Without Gate” condition, both image-level and feature-level perturbations contribute to performance enhancement, and their combined utilization further improves performance. ❷ A comparison between “Feature + Image” and “Ours(FP(d1))” reveals that incorporating the gate mechanism helps distinguish challenging samples, reducing training difficulty and improving performance. ❸ Under the “With Gate” condition, contrasting “FP(d4)”, “FP(d1,d4)”, and “Ours(FP(d1))” indicates that disturbing deep-level features can introduce excessive disruptions, increasing training complexity. Disturbing only the shallow-level features appears more robust than disturbing the deep-level features. ❹ Under the “With Gate” condition, “Ours(FP(d1))” employs ResNet50 as the backbone. In comparison with the approach based on “ResNet101”, it is observed that with a limited number of labels, the performance of the ResNet50-based is slightly better than that of the ResNet101-based. As the number of labels increases, the performance of ResNet101-based surpasses that of the ResNet50-based.
TABLE IV: Ablation study on WHU-CD dataset.
Method 5% 10% 20% 40%
IoU OA IoU OA IoU OA IoU OA
Without Gate Sup-only 41.4 96.36 55.0 97.28 49.3 96.98 65.6 98.26
Feature 68.0 98.49 73.9 98.80 74.3 98.77 82.0 99.21
Image 77.4 99.03 80.0 99.14 77.5 98.97 82.2 99.22
Feature + Image 81.7 99.24 81.6 99.21 79.2 99.07 84.8 99.35
With Gate FP(d4) 79.4 99.13 78.0 99.03 82.9 99.25 83.4 99.28
FP(d1, d4) 78.5 99.08 81.3 99.20 79.4 99.07 84.4 99.33
Ours(FP(d1)) 83.0 99.30 82.8 99.28 84.2 99.31 85.4 99.37
ResNet101 82.6 99.28 82.3 99.24 82.1 99.21 85.7 99.39
Sensitivity of the gate mechanism. We apply the optimal threshold on WHU-CD due to space limitations. The experimental results are shown in Fig 4. High ratio lead to unstable training as more samples are perturbed, while low ratio may miss perturbation opportunities. Median threshold strikes a balance.
Figure 4: Performance comparison with different perturbation ratios under 5% labeled training data on WHU-CD dataset.
IV Conclusion
In this work, we introduce a novel gate-guided two-level perturbation consistency regularization-based SSCD method. This approach ensures image-level strong-to-weak consistency and feature-level perturbation consistency, enabling efficient utilization of a large amount of unlabeled data. Moreover, we design a gate module to intelligently access the need of feature perturbations. Extensive experiments on six benchmark datasets validate the effectiveness and superiority of our proposed method. In the future, we will further explore enabling the network to autonomously determine the most effective perturbation strategies tailored to various data scenarios.